# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v2

@@ H.1 Estimates of the Matrix Elements and the Frobenius and Operator norms in the Localized Phase
$\left|\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}\right|\sim O\left(\frac{1}{\sqrt{\xi}}\right)^{2}\times O\left(\sqrt{\xi}\right)\;\sim\;O\left(\frac{1}{\sqrt{\xi}}\right)\leavevmode\nobreak\ .$ (87)
Hence, for the squared Frobenius norm of the charge transport capacity operator $\hat{M}_{i_{0}}^{(0)}$, which up to a factor of 1/41/41 / 4 coincides with the standard squared Frobenius norm of the matrix $M_{i_{0}}^{(0)}$ (cf. Appendix B), we obtain
$\big{\|}\hat{M}_{i_{0}}^{(0)}\big{\|}_{F}^{2}\sim O\left(\xi\times\xi\right)\times O\left(\frac{1}{\sqrt{\xi}}\right)^{2}\sim O(\xi)\leavevmode\nobreak\ .$ (88)
@@
$\big{\|}\hat{M}_{i_{0}}^{(0)}\big{\|}_{F}\sim O(\sqrt{L})\leavevmode\nobreak\ ,\qquad\big{\|}\hat{M}_{i_{0}}^{(0)}\big{\|}_{\text{op}}\sim O(L)\leavevmode\nobreak\ .$ (90)
@@ H.2 Concise Expression for the Frobenius Norm and More Accurate Estimates
$\|\hat{M}_{i_{0}}\|_{F}^{2}$ $=\frac{1}{4}\sum_{\alpha\neq\beta}\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}^{2}\leavevmode\nobreak\ ,$ (91)
@@
$=i_{0}-\sum_{\alpha}\left(\sum_{j=1}^{i_{0}}\phi_{j\alpha}^{2}\right)^{2}\leavevmode\nobreak\ ,$
@@ Appendix I Operator Norm Data for $\hat{M}^{(0)}$
Changing the wavenumber, however, does shift which lattice sizes happen to be very close to commensurate with the underlying potential. We note the pronounced peaks at $L=40$ and $L=70$, where the $L=70$ peak is absent from the $k=\sqrt{2}$ case. These peaks can be attributed to the same mechanism discussed in Sec. III.4.2. Specifically, $(L+1)\times k=41\times\sqrt{3}\approx 71$ and $71\times\sqrt{3}\approx 123$, both of which are close to integers. This supports our interpretation that such peaks arise from near-commensurability of the quasiperiodic potential at these specific system sizes.

### v3

@@ H.1 Estimates of the Matrix Elements and the Frobenius and Operator norms in the Localized Phase
$\left|\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}\right|\sim O\left(\frac{1}{\sqrt{\xi}}\right)^{2}\times O\left(\sqrt{\xi}\right)\;\sim\;O\left(\frac{1}{\sqrt{\xi}}\right)\penalty 10000\ .$ (90)
Hence, for the squared Frobenius norm of the charge transport capacity operator $\hat{M}_{i_{0}}^{(0)}$, which up to a factor of $1/4$ coincides with the standard squared Frobenius norm of the matrix $M_{i_{0}}^{(0)}$ (cf. Appendix B), we obtain
$\big\|\hat{M}_{i_{0}}^{(0)}\big\|_{F}^{2}\sim O\left(\xi\times\xi\right)\times O\left(\frac{1}{\sqrt{\xi}}\right)^{2}\sim O(\xi)\penalty 10000\ .$ (91)
@@
$\big\|\hat{M}_{i_{0}}^{(0)}\big\|_{F}\sim O(\sqrt{L})\penalty 10000\ ,\qquad\big\|\hat{M}_{i_{0}}^{(0)}\big\|_{\text{op}}\sim O(L)\penalty 10000\ .$ (93)
@@ H.2 Concise Expression for the Frobenius Norm and More Accurate Estimates
$\|\hat{M}_{i_{0}}\|_{F}^{2}$ $=\frac{1}{4}\sum_{\alpha\neq\beta}\left(M_{i_{0}}^{(0)}\right)_{\alpha\beta}^{2}\penalty 10000\ ,$ (94)
@@
$=i_{0}-\sum_{\alpha}\left(\sum_{j=1}^{i_{0}}\phi_{j\alpha}^{2}\right)^{2}\penalty 10000\ ,$
@@ Appendix I Operator Norm Data for $\hat{M}^{(0)}$
Changing the wavenumber, however, does shift which lattice sizes happen to be very close to commensurate with the underlying potential. We note the pronounced peaks at $L=40$ and $L=70$, where the $L=70$ peak is absent from the $k=\sqrt{2}$ case. These peaks can be attributed to the same mechanism discussed in Sec. III.4.3. Specifically, $(L+1)\times k=41\times\sqrt{3}\approx 71$ and $71\times\sqrt{3}\approx 123$, both of which are close to integers. This supports our interpretation that such peaks arise from near-commensurability of the quasiperiodic potential at these specific system sizes.
@@
Appendix K Finite-Size Scaling in the Quasiperiodic Model
Figure 25: Same data as in Fig. 6(e) in the localized regime $h>2$, but with the system size on the x-axis scaled by the localization length $\xi=1/\log(h/2)$, and the y-axis scaled by a phenomenological scaling of $\xi^{-3}$.
Figure 26: Median of $\|\hat{M}^{(0)}\|_{F}$ in the quasiperiodic model, with the x-axis scaled by the localization length $\xi=1/\log(h/2)$, and the y-axis scaled roughly by $\sqrt{\xi}$ (which is a phenomenologically expected scaling). The data is the same as in Fig. 12(e), showing only the localized regime $h>2$, with more added $h$ values for more thorough scaling collapse analysis, which is easy to do in the non-interacting case. The collapse of the curves suggests the finite-size scaling $\mathrm{med}(\|\hat{M}^{(0)}\|_{F})\sim\sqrt{\xi}f(L/\xi)$.
Figure 27: Same data of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as in Fig. 16(e), but with the x-axis scaled by the localization length $\xi=1/\log(h/2)$, showing data only in the localized regime $h>2$ of the non-interacting starting model. The y-axis is scaled by a phenomenological scaling $\xi^{4.4}$.
In this section, we briefly show a finite-size scaling analysis for quantities relating to the quasiperiodic Aubry-André model, specifically in the localized regime, similarly to the analysis in the Anderson model in the main text. The analysis is easier in the Aubry-André model, as it admits an exact formula for the localization length $\xi$ [27]:
$\xi=1/\log(2/h).$ (96)
First, we look at the LIOM corrections in the quasiperiodic model, and wish to collapse the $\mathrm{med}\|\tilde{n}^{(1)}\|_{F}$ curves, just like in the case with the LIOM corrections of the Anderson model. We scale the x-axis by $\xi$, and phenomenologically scale the y-axis by a power of $\xi$. The results of this collapse are shown in Fig. 25.
We next consider the noninteracting charge transport capacity $\|\hat{M}^{(0)}\|_{F}$. Inspired by the case for the Anderson model, a natural finite-size scaling ansatz to assume is:
$\mathrm{med}\|\hat{M}^{(0)}\|_{F}\sim\sqrt{\xi}f(L/\xi).$ (97)
We show the collapse of the curves using this ansatz in Fig. 26, supporting the ansatz and illustrating the finite-size effects.
Finally, we study the interaction correction to the charge transport capacity $\hat{M}^{(1)}$ in the localized phase While we do not have as much data as for $\hat{M}^{(0)}$, we can still attempt to collapse the $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ curves, just like in the case with the charge transport capacity corrections of the Anderson model– see Fig. 27 for the results.